# CoverUp: Coverage-Guided LLM-Based Test Generation Juan Altmayer Pizzorno

### v1

@@ III Technique
without any changes to the code under test [27, 28].
@@
Figure 8: Example of a flaky test:
@@
By repeating the test a few times, CoverUp is likely to find
@@
While CoverUp does not attempt to detect flaky tests by default, it
supports passing custom arguments to pytest.
This approach lets the user specify that each test should be repeated a number
of times using the pytest-repeat plugin.
Flaky tests are then more likely to fail, leading CoverUp to give the
LLM a chance to correct the problem.
III-FOther Technical Challenges
@@
Even though CoverUp limits the amount of time spent waiting on responses and
test executions, if each code segment was processed serially, CoverUp would
take an unacceptable amount of time creating tests for packages of nontrivial size.
CoverUp instead prompts for tests and verifies them asynchronously, using the Python
@@
As a result, during CoverUp’s evaluation (Section IV) we observe
@@
CoverUp handles a number of timeout and other error conditions automatically.
To avoid exceeding the rate limits imposed by OpenAI, CoverUp spreads out its
requests using a leaky bucket [29] scheme implemented by the
@@
Additionally, CoverUp can create and use checkpoint files, allowing the user
@@ IV Evaluation
RQ1:
How does CoverUp’s coverage compare to the previous state of the art?
(Section IV-B)
RQ2:
How does CoverUp’s coverage compare to CodaMosa using a state-of-the-art LLM?
(Section IV-C)
RQ3:
How effective is CoverUp’s continued dialogue at increasing coverage?
(Section IV-D)
IV-AExperimental Setup
@@
We utilize a benchmark suite collated by the authors
of CodaMosa [2], available at https://github.com/microsoft/codamosa;
we refer to it as suite CM.
It stems originally from 35 projects used in the evaluation of
BugsInPy [30] and
Pynguin [31].
The suite is made up of a subset of the Python modules in these projects, selected
to focus on cases where Pynguin has difficulty achieving full coverage:
after removing the modules that failed to produce results, removing the modules
where Pynguin achieved full coverage within one minute, and downsampling the
number of modules with a common parent, the suite contains 486 benchmark modules
from 27 projects.
Given suite CM’s focus, evaluating CoverUp only on it would leave open
the question of how it performs on code where SBST, or more precisely, the
original pre-CodaMosa Pynguin already performed well.
For that reason, to investigate RQ1 we create an additional suite from the same source,
using only modules for which Pynguin achieved full coverage, naming it PY.
Further, if a module is already in CM, indicating that Pynguin failed at least once
to achieve full coverage, we exclude it from PY.
By evaluating CoverUp on both suites, we obtain a fuller
picture of CoverUp’s performance.
@@
We use two versions of CodaMosa, the previous state of the art in
automated case test generation, as baselines for evaluation on the CM suite:
the original version that uses the Codex model (“CodaMosa (codex)”) and
our adapted version that uses the same model as CoverUp (“CodaMosa (gpt4)”).
To investigate RQ1, we utilize the tests CodaMosa (codex) originally generated
using its best performing configuration (codamosa-0.8-uninterp),
available from its “dataset” package at https://github.com/microsoft/codamosa-dataset.
The dataset includes 16 runs of CodaMosa for each of the benchmark modules.
Rather than use Pynguin’s coverage measurements from when the tests were generated,
which differs subtly from SlipCover in how it counts branch coverage, as we
describe below,
we re-execute its tests and measure the resulting coverage (Section IV-A).
For RQ2, we adapt CodaMosa to use the same LLM as CoverUp.
The code-davinci-002 (“codex”) LLM used by the original CodaMosa,
which is no longer available, used code completion prompts, where
the LLM, given a code fragment, responds with a suggestion on how to complete it.
CodaMosa (codex) sent that LLM portions of the code under test to provide
it with context as well as the header of a unit test function for that code.
We create CodaMosa (gpt-4) by modifying it to use OpenAI’s chat API and
to insert instructions requesting a code completion before CodaMosa (codex)’s
original prompt.
Figure 10 shows an example of the resulting prompt.
We configure CodaMosa (gpt4) to use OpenAI’s gpt-4-1106-preview LLM,
setting a temperature of 0, the same employed by CoverUp in this evaluation, but
otherwise leave the configuration identical to that of codamosa-0.8-uninterp.
Figure 9 shows the details.
Figure 9: Configuration for CodaMosa (gpt4):
This configuration selects the same LLM and temperature as
CoverUp, but is otherwise based on CodaMosa’s
best performing configuration, codamosa-0.8-uninterp.
12Figure 10: Example prompt after adapting CodaMosa to a chat LLM:
The figure shows a CodaMosa completion prompt, adapted for use
with a chat LLM.
The prompt prefixes instructions to complete the code (1) before
including the original prompt (2).
The code shown is from the flutils package.
@@
We measure all coverage using SlipCover [22], including
those for CodaMosa (codex) generated tests.
We do not use Pynguin’s coverage measurements from when CodaMosa (codex)
generated them because Pynguin measures branch coverage differently from
both SlipCover and Python’s de facto standard coverage tool, coverage.py [32].
Before executing a program, the Python interpreter compiles the source code to bytecode.
Pynguin counts the jump instructions used to implement conditionals as branches and thus
measures clause coverage [33, 22].
Figure 11 shows an example with a two-clause conditional
and the resulting bytecode.
Counting such jump instructions as branches increases the total number of possible
“branches” and can thus lead to both inflated and understated coverage percentual
numbers, depending on how many of those were followed during execution.
Figure 11: Python bytecode for two-clause conditional:
Python compiles the two-clause conditional if A and B into
bytecode that includes two jump instructions.
If a coverage tool counts these as branches,
it is really measuring clause coverage.
@@
We execute both the CodaMosa tests and CoverUp using the codamosa-docker
@@
Unfortunately, pip fails to install various of these requirements.
@@
GPT-generated tests often assume that certain Python modules are available on the system.
Before each CoverUp run, we install the modules mock, pytest-datadir,
pytest-mock, responses and hypothesis.
CoverUp supports automatically installing modules required by tests, but our evaluation
does not use that functionality.
@@
We configure CoverUp to use OpenAI’s recent gpt-4-1106-preview LLM, setting its “temperature” to zero,
and do not limit the number of output tokens.
We leave CoverUp’s target code segment size (see Section III-A)
at its default of 50 lines, and do not automatically repeat test executions
to look for flaky tests (see Section III-E).
@@
We utilize CoverUp’s disable failing test resolution method that
simply disables any tests that fail during integration check
(see Section III-D), and do not pass any extra
pytest flags (see Section III-E).
IV-BComparison to the previous state of the art
@@
suite, using CodaMosa (codex) as the baseline.
Figure 12 shows the line, branch and combined line and
branch coverage obtained by CoverUp and the baseline over the entire benchmark suite;
Figure 13 shows the median of those measurements
on a module-by-module basis.
@@
line and branch coverage than CodaMosa (codex) both measuring over the entire
@@
Figure 12: [RQ1, RQ2] CoverUp yields higher overall coverage:CoverUp yields higher line, branch and line+branch coverage
over the entire benchmark source than both CodaMosa (codex),
and CodaMosa (gpt4).
Figure 13: [RQ1, RQ2] CoverUp yields higher median per-module coverage:CoverUp yields higher median line, branch and line+branch coverage
on a module-by-module basis than than both CodaMosa (codex) and CodaMosa (gpt4).
Figure 14: [RQ1] CoverUp coverage on the PY suite:CoverUp yields near 100% overall and 100% median coverage
@@
Figure 15: [RQ3] Contribution of iterative coverage-guided approach:
almost half (42.9%) of the successes were obtained
by iterating and providing updated coverage information, demonstrating its importance.
Figure 16: [RQ1] CoverUp yields higher coverage than the state of the art:
CoverUp and CodaMosa (codex) for the various benchmark modules in the CM benchmark suite.
@@
Given CoverUp’s near 100% overall coverage and 100% median per-module coverage
@@
[RQ1] Summary:CoverUp achieves significantly higher coverage than CodaMosa (codex),
the previous state of the art.
IV-CHow does CoverUp’s coverage compare to CodaMosa using a state-of-the-art LLM?
To investigate RQ2, we run CodaMosa (gpt4) on the CM suite and measure the coverage it obtains.
Figures 12 and 13 show the results,
measuring coverage over the entire benchmark suite and on a per-module basis.
As the figures show, CoverUp also outperforms CodaMosa (gpt4), achieving higher coverage
both across the entire suite and a per-module basis.
We also observe that CodaMosa (gpt4)’s performance differs little from that of
CodaMosa (codex), with overall coverage and per-module line and combined line and
branch measurements within 2% of each other.
[RQ2] Summary:CoverUp still outperforms CodaMosa using a state-of-the-art LLM.
IV-DHow effective are CoverUp’s continued dialogue?
@@
was generated immediately upon the first prompt, after continuing the chat with a second prompt,
or after a third prompt.
By default, and also in our evaluation, CoverUp continues the conversation for at most
two additional prompts.
@@
As the figure shows, while the success rate decreases with each iteration, almost
half of the successes (49.2%) were obtained through iterative refinement of the prompt,
@@
[RQ3] Summary: continuing the chat contributes to nearly half of successes,
@@ V Threats to Validity
We utilize the same benchmark suite as CodaMosa to facilitate
a direct comparison to the state of the art.
@@
CodaMosa’s original evaluation environment failed to install various
Python modules that are prerequisites to the applications used as benchmarks.
@@
For CoverUp, the modules contributed by the ansible package are
the most affected, with 326 test generation attempts being aborted due to
missing modules.
@@
All of CoverUp’s development and evaluation was performed using
OpenAI GPT-4 and GPT-4 Turbo models.
@@
are independent of the LLM, its ultimate performance naturally
depends heavily on the model’s ability to interpret its prompts
and generate tests as requested.
VI Discussion and Future Work
In an attempt to guide the LLM towards useful tests, CoverUp requests
that tests contain assertions in its initial prompt (Section III-B)
and also treats as an error any tests that fail (Section III-C).
It is still possible, however, that the assertions generated by the LLM
are not sufficient to verify the test postconditions.
Even though this paper focuses on case generation and coverage,
rather than the test oracle, we examined a sample of
the assertions in the generated tests.
Many of the assertions seemed appropriate, verifying that the code under test had
the intended effect, but others, for example, merely checked that the
state had changed.
In future work, we intend to extend CoverUp to evaluate the assertions,
possibly employing mutation testing [34, 10],
and contining the dialogue (Section III-C) to request improvements.
Another case that is likely to be a good candidate for continued prompting is
when the LLM generates a test requiring Python modules that are not present
on the system.
As we describe in Section III-C, CoverUp detects this situation
and optionally attempts to install or records a list of missing modules for
later installation; it could, however, instead prompt the LLM to adjust or
re-formulate the test so that the module is not necessary.
Initial experiments suggest that this may be a viable way to automatically
resolve such cases.
A number of initial prompt test generation failures could be attributed
to CoverUp not providing enough context in its prompts, such as related
definitions from other source files.
In future work, we intend to experiment with ways to provide additional
context.
In particular, OpenAI’s function calling feature, which allows
a model to request additional information in response to a prompt, may
work well for this purpose, reducing the need for continued prompting.
Finally, we developed and evaluated CoverUp exclusively using OpenAI
models (GPT-4, GPT-4 Turbo); we intend to adapt CoverUp to work with
other models as well.
We believe that adapting it for use with on-premises LLMs might be of
special interest to industry practitioners, as it would help protect
intellectual property.
VII Conclusion
@@
We show that coupling coverage information with prompting, and interatively
@@
test case generator
@@
CodaMosa (codex)
CodaMosa (gpt4)
TABLE I: [RQ1, RQ2] Overall coverage obtained on the CM suite.
test case generator
line
branch
line + branch
CoverUp
CodaMosa (codex)
CodaMosa (gpt4)
TABLE II: [RQ1, RQ2] Median per-module coverage obtained on the CM suite.
@@
line + branch
overall coverage
per-module median
TABLE III: [RQ1] CoverUp coverage on the PY suite.
@@
TABLE IV: [RQ3] Contribution of iterative coverage-guided approach to successes.
[1]
S. Lukasczyk and G. Fraser, “Pynguin: automated unit test generation for
python,” in Proceedings of the ACM/IEEE 44th International Conference
on Software Engineering: Companion Proceedings, ser. ICSE ’22. ACM, May 2022. [Online]. Available:
http://dx.doi.org/10.1145/3510454.3516829
[2]
C. Lemieux, J. P. Inala, S. K. Lahiri, and S. Sen, “Codamosa: Escaping
coverage plateaus in test generation with pre-trained large language
models,” in 2023 IEEE/ACM 45th International Conference on Software
Engineering (ICSE), 2023, pp. 919–931.
[3]
C. Boyapati, S. Khurshid, and D. Marinov, “Korat: automated testing based on
java predicates,” in Proceedings of the 2002 ACM SIGSOFT International
Symposium on Software Testing and Analysis, ser. ISSTA ’02. New York, NY, USA: Association for Computing
Machinery, 2002, p. 123–133. [Online]. Available:
https://doi.org/10.1145/566172.566191
[4]
B. P. Miller, L. Fredriksen, and B. So, “An empirical study of the reliability
of UNIX utilities,” Commun. ACM, vol. 33, no. 12, p. 32–44, dec
1990. [Online]. Available: https://doi.org/10.1145/96267.96279
[5]
A. Fioraldi, A. Mantovani, D. Maier, and D. Balzarotti, “Dissecting American
Fuzzy Lop: A fuzzbench evaluation,” ACM Trans. Softw. Eng.
Methodol., vol. 32, no. 2, mar 2023. [Online]. Available:
https://doi.org/10.1145/3580596
[6]
C. Pacheco, S. K. Lahiri, M. D. Ernst, and T. Ball, “Feedback-directed random
test generation,” in ICSE 2007, Proceedings of the 29th International
Conference on Software Engineering, Minneapolis, MN, USA, May 2007, pp.
75–84.
[7]
P. Godefroid, N. Klarlund, and K. Sen, “Dart: directed automated random
testing,” in Proceedings of the 2005 ACM SIGPLAN Conference on
Programming Language Design and Implementation, ser. PLDI ’05. New York, NY, USA: Association for Computing
Machinery, 2005, p. 213–223. [Online]. Available:

### v2

@@ 3.6. Handling Flaky Tests
without any changes to the code under test (Parry et al., 2021; Lam et al., 2020; Gruber et al., 2024).
@@
Figure 9. Example of a flaky test:
@@
CoverUp executes each test a few times, making it more likely to find
@@
Our implementation uses the pytest-repeat module to execute each
newly generated test multiple times, making any flaky tests more likely to fail.
If a test fails, CoverUp continues the chat, giving the LLM a chance
to correct the problem.
3.7. Other Technical Challenges
@@
Even though our implementation limits the amount of time spent waiting on responses and
test executions, if each code segment was processed serially, it would
take an unacceptable amount of time to create tests for packages of nontrivial size.
Instead, it prompts for tests and verifies them asynchronously, using the Python
@@
As a result, during CoverUp’s evaluation (Section 4) we observe
@@
Our implementation handles a number of timeout and other error conditions automatically.
To avoid exceeding the rate limits imposed by OpenAI, it spreads out its
requests using a leaky bucket scheme (Turner, 1986) implemented by the
@@
Additionally, we implement checkpointing to files, allowing the user
@@ 4. Evaluation
RQ1::
Does the coverage of CoverUp’s generated tests improve upon the state of the art?
(Section 4.2)
RQ2::
How effective is CoverUp compared to simply prompting an LLM for tests?
(Section 4.3)
RQ3::
How effective are CoverUp’s continued dialogues at increasing coverage?
(Section 4.4)
4.1. Experimental Setup
@@
We utilize three benchmark suites:
CM, a benchmark suite on which Pynguin struggles to obtain high coverage.
Originally collated by the authors of CodaMosa (Lemieux et al., 2023) and available
from https://github.com/microsoft/codamosa, it is derived from 35 open source
projects used in the evaluation of BugsInPy (Widyasari et al., 2020) and
Pynguin (Lukasczyk et al., 2023).
It contains 100,000 lines of code across 425 Python modules.
PY, a set of modules originally excluded from CodaMosa’s suite because
Pynguin already performs well on it.
We evaluate CoverUp on these modules so as not to leave open the question of
how well it performs on code where Pynguin / SBST already performed well.
It contains 5,000 lines of code across 84 Python modules.
MT, a dataset of functions extracted from the HumanEval dataset (Chen et al., 2021),
originally used in the evaluation of MuTAP (Dakhel et al., 2024).
We use this dataset to enable comparisons to MuTAP, since its implementation
lacks the support needed to to run on arbitrary Python packages.
It contains 2,000 lines of code across 163 functions.
CM excludes the mimesis, sanic and thef*ck packages
from the original CodaMosa suite: mimesis is a package for generating random
data, making it difficult to generate non-flaky tests for it.
mimesis was included in CodaMosa’s original evaluation, but the tests CodaMosa
generates do not contain any assertions, and thus do not fail in the face of randomly generated values.
The sanic and thef*ck packages require modules that either are no longer
available or conflict with CoverUp.
@@
To examine RQ1, we compare against two versions of CodaMosa and
four versions of MuTAP, varying their LLM and prompt type:
CodaMosa (codex), the original version that uses the Codex LLM.
Since Codex is no longer available, we use original tests from CodaMosa’s
evaluation, using its best performing configuration (0.8-uninterp),
available from https://github.com/microsoft/codamosa-dataset;
CodaMosa (gpt4o), our adapted version that uses the same model as
CoverUp, to help rule out differences in performance due to the use of
different LLMs.
MuTAP (codex few-shot), the original version that uses the Codex LLM,
using a “few-shot” prompt, where two examples of test generation are included
with the prompt.
Since Codex is no longer available, we use original tests from MuTAP’s
evaluation, available from https://github.com/ExpertiseModel/MuTAP;
MuTAP (codex zero-shot), also the original Codex version, but using
a prompt without examples.
Here, too, we use original tests from MuTAP’s evaluation;
MuTAP (gpt4o few-shot), our adapted version that uses the same model
as CoverUp, with newly generated tests using a “few-shot” prompt;
MuTAP (gpt4o zero-shot), our adapted version, but using a “zero-shot”
prompt.
To create the gpt4o versions of CodaMosa and MuTAP, we modify these to use
OpenAI’s chat API and, in the case of CodaMosa, insert instructions requesting
a code completion before its original code completion prompt.
For MuTAP, we insert a prompt requesting that it respond in the same format as
a Codex model.
We were unable to empirically evaluate (Ryan et al., 2024) as at the time of writing
it remains unavailable: its authors have indicated that it is not yet publicly available.
For RQ2, we create an “CoverUp (ablated)”, a version of CoverUp which utilizes
a nearly identical prompt, except in that it does not specify how the code lacks
coverage or tags lines lacking coverage, does not include computed import
statements, does not offer a tool function for additional context, and does not
continue the chat in case of errors or lack of coverage.
@@
We run all generated tests in isolation (see Section 3.5)
and measure all coverage using SlipCover (Altmayer Pizzorno and Berger, 2023).
@@
We evaluate both CodaMosa and CoverUp using the codamosa-docker
@@
Unfortunately, pip fails to install a number of these requirements.
@@
For MuTAP and the tests it generates, we use Python 3.9.12: MuTAP uses MutPy (Hałas, 2019)
for mutation testing, which requires older versions of Python.
@@
We configure CoverUp to use OpenAI’s gpt-4o-2024-05-13 LLM,
setting its “temperature” to zero, and do not limit the number of output tokens.
We leave CoverUp’s target code segment size (see Section 3.1)
at its default of 50 lines, and automatically repeat test executions
to look for flaky tests (see Section 3.6).
@@
Since GPT-generated tests often assume the presence of certain Python modules,
we configure CoverUp to automatically install any such modules if missing.
4.2. Comparison to the previous state of the art
@@
suite, using CodaMosa (codex) and CodaMosa (gpt4o) as baselines.
Figure 10 shows, on the left, the line, branch and combined line and
branch coverage obtained by CoverUp and the baselines over the entire benchmark suite;
on the right, it shows the median of those measurements on a module-by-module basis.
@@
line and branch coverage than both CodaMosa baselines, both measuring over the entire
@@
Across the entire benchmark suite, CoverUp achieves 64% (vs. 54% and 51%) line coverage,
49% (vs. 34% and 27%) branch coverage, and 60% (vs. 49% and 45%) line+branch coverage;
on a per-module basis, CoverUp achieves 82% (vs. 63% and 54%) line coverage,
67% (vs. 36% and 19%) branch coverage, and 80% (vs. 56% and 47%) line+branch coverage.
We also observe that CodaMosa (gpt4o)’s performance falls slightly behind that of
CodaMosa (codex), with overall coverage measurements within 5% of each other.
As this difference shows, a newer model does not necessarily lead to higher performance.
We next compare CoverUp to MuTAP, evaluating it on the MT suite.
Figure 11 shows, on the left, the line, branch and combined line and
branch coverage obtained by CoverUp and the baselines over the entire benchmark suite;
on the right, it shows the median of those measurements on a module-by-module basis.
As the figures show, CoverUp achieves higher line, branch and combined line and branch
coverage than the MuTAP baselines when measuring over the entire benchmark,
and greater or equal coverage when measuring on a per-module basis.
Across the entire benchmark suite, CoverUp achieves 90% (vs. 82% to 16%) line coverage,
89% (vs. 73% to 0%) branch coverage, and 90% (vs. 79% to 10%) line+branch coverage;
on a per-module basis, like its baselines, CoverUp achieves 100% on all metrics,
except for MuTAP (gtp4o few-shot), which achieves extremely low coverage
(17%, 0% and 11%).
We investigate MuTAP (gpt4o few-shot)’s extremely low performance and discover that
its original few-shot prompt confuses GPT-4o, leading it to generate tests for
the examples that make up the “few shot”, rather than for the function under test.
When this happens, even if the resulting tests run to completion, they do not
contribute to coverage.
As the consistently high per-module median coverage values indicate, the MT suite does not
contain particularly challenging code.
In fact, the functions in the suite are entirely self contained, only rarely utilize external code,
and when they do so, from standard libraries, and commonly include type annotations.
All of these characteristics greatly simplify the task of test generation.
By contrast, the functions in the CM suite have numerous dependencies, often use external
code and are generally unannotated.
Figure 10. [RQ1] CoverUp yields higher overall and median per-module coverage:
Across the board, CoverUp yields higher coverage than CodaMosa,
whether measured over the entire suite or on a module-by-module basis.
Figure 11. [RQ1] CoverUp yields higher overall and median per-module coverage:CoverUp yields higher or equal coverage than MuTAP,
whether measured over the entire suite or on a module-by-module basis.
Figure 12. [RQ1] CoverUp yields higher coverage than the state of the art:
CoverUp and CodaMosa for the various benchmark modules in the CM benchmark suite,
sorted by percentage of coverage increase.
Figure 13. [RQ1] CoverUp also performs well on the PY suite:CoverUp yields 100% median and near 100% overall coverage
@@
Figure 14. [RQ3] Contribution of iterative coverage-guided approach:
nearly 40% of the successes were obtained
by iterating and providing coverage and error feedback,
demonstrating its importance.
@@
Given CoverUp’s 100% median per-module coverage and near 100% overall coverage
@@
[RQ1] Summary:CoverUp achieves significantly higher coverage than
both CodaMosa and MuTAP, outperforming the state of the art.
4.3. How effective is CoverUp compared to simply prompting an LLM for tests?
Given many LLMs’ near-human performance on various tasks, including software
engineering tasks (Bubeck et al., 2023), it is reasonable to ask just how
much CoverUp contributes to performance.
To address this question, we compare CoverUp to an ablated LLM-only
version that utilizes a nearly identical prompt, but which lacks all other
CoverUp functionality (coverage information, precomputed computed
import statements, tool function for additional context
and continued chats).
As Figure 15 shows,
CoverUp achieves substantially higher line, branch and combined line and branch coverage
than CoverUp (ablated) both measuring over the entire benchmark code base
and on a per-module basis.
Across the entire benchmark suite, it achieves 64% (vs. 39%) line coverage,
49% (vs. 19%) branch coverage, and 60% (vs. 34%) line+branch coverage;
on a per-module basis, it achieves 82% (vs. 44%) line coverage,
67% (vs. 7%) branch coverage, and 80% (vs. 39%) line+branch coverage.
Figure 15. [RQ2] CoverUp contributes significantly to overall performance:
Across the board, CoverUp yields higher coverage in comparison to
the ablated (LLM only) version, whether measured over the entire suite
or on a module-by-module basis.
[RQ2] Summary:CoverUp contributes significantly to overall performance;
its performance is not just due to the LLM employed.
4.4. How effective are CoverUp’s continued dialogues?
@@
was generated immediately upon the first prompt or after continuing the chat with a second
or third prompt
(by default, and also in our evaluation, CoverUp continues the conversation for at most
two additional prompts).
@@
As the figure shows, while the success rate decreases with each iteration, about
40% of the successes were obtained through iterative refinement of the prompt,
@@
[RQ3] Summary: continuing the chat contributes almost half its successes,
@@ 5. Threats to Validity
We utilize the CM and PY benchmark suites to evaluate CoverUp’s
performance on both challenging and less challenging code, as well
as to facilitate including Codex-based results;
we utilize the MT benchmark suite because it is supported by MuTAP
and it, too, facilitates including Codex-based results.
@@
CodaMosa’s original evaluation environment failed to install a number of
Python modules that are prerequisites for the applications used as benchmarks.
@@
CoverUp was developed and evaluated using OpenAI’s GPT-4, GPT-4 Turbo
and GPT-4o models.
@@
are independent of the LLM
and, as Section 4.3 shows, CoverUp significantly
outperforms an ablated version of itself that just relies on the LLM’s capabilities,
its ultimate performance naturally depends on the model’s ability to
interpret its prompts and generate tests as requested.
6. Conclusion
@@
We show that coupling coverage information with prompting, and iteratively
@@
Table 1. [RQ1, RQ2] CoverUp outperforms CodaMosa on the CM suite and
far outperforms itself when ablated to simply rely on LLM performance.
overall coverage
median per-module coverage
test generator
line
branch
line + branch
@@
MuTAP (codex few-shot)
MuTAP (codex zero-shot)
MuTAP (gpt4o zero-shot)
MuTAP (gpt4o few-shot)
Table 2. [RQ1] CoverUp outperforms MuTAP’s coverage on the MT suite.
@@
l+b
module median
overall
Table 3. [RQ1] CoverUp’s performance is not limited to “challenging” code: results for PY suite
@@
Table 4. [RQ3] Contribution of iterative coverage-guided approach to successes.
A Large-Scale Evaluation of Automated Unit Test
Generation Using EvoSuite.
ACM Trans. Softw. Eng. Methodol.
24, 2, Article 8
(December 2014),
42 pages.
https://doi.org/10.1145/2685612
Fraser and Zeller (2012)
Gordon Fraser and
Andreas Zeller. 2012.
Mutation-Driven Generation of Unit Tests and
Oracles.
IEEE Transactions on Software Engineering
38, 2 (2012),
278–292.
https://doi.org/10.1109/TSE.2011.93
Godefroid et al. (2005)
Patrice Godefroid, Nils
Klarlund, and Koushik Sen.
2005.
DART: directed automated random testing. In
Proceedings of the 2005 ACM SIGPLAN Conference on
Programming Language Design and Implementation (Chicago, IL, USA)
(PLDI ’05). Association for
Computing Machinery, New York, NY, USA,
213–223.